# Kekulé Superconductivity in Twisted Magic Angle Bilayer Graphene

### v1

@@ Appendix B U/V-transitions in the DOS
As discussed in the main text, the U/V transition can be understood through the single-band analysis presented in Eq. (28). For a typical order parameter, say $\Delta\sim 1$ meV, the associated Bogoliubov Fermi surface is small, as depicted in Fig. 6. This originates from the condition $\delta\epsilon/W\ll 1$, which indicates that the inversion symmetry breaking is quite weak. Consequently the “V” shape has a rather sharp minimum.
@@
Additionally, the energy scale of the coherence-peaks, $\Delta_{t}$, can be extracted from the DOS. We find that it is smaller than the order parameter, and satisfies $\Delta_{t}/\Delta\approx 0.4<1$. Noteably, although the ratio $2\Delta_{t}/T_{c}$ is smaller than $2\Delta/T_{c}$, it remains significantly larger than the weak-coupling BCS value.
@@ Appendix E Acknowledgement
We thank Kevin Nuckolls, Ivar Martin, Bitan Roy, Zhiqiang Wang and Ananth Malladi for helpful discussions. We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.

### v2

@@ Appendix B U/V-transitions in the DOS
As discussed in the main text, the U/V transition can be understood through the analysis presented in Eq. (10). For a typical order parameter, say $\Delta\sim 0.3$ meV, the associated Bogoliubov Fermi surface is small, as depicted in Fig. 5. This implies that the “V” shape has a rather sharp minimum.
@@
In Fig. 8 we plot the effective zero bias conductance as well as the order parameter for a strong attraction case where the tunneling curve assumes a U-shape. This represents a prediction as, thus far, there are no published curves for this zero bias conductance. In contrast to Fig. 6 at weaker attraction, here, as temperature increases the zero bias density of states remains gapped until crossing over to the behavior associated with the V-shaped regime where a Bogoliubov Fermi surface is present.
@@ Appendix E Acknowledgement
We thank Kevin Nuckolls, Hyunjin Kim, Ivar Martin, Bitan Roy, Stevan Nadj-Perge, Senthil Todadri, Zhiqiang Wang, Ananth Malladi for helpful discussions. We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.